OpenStudy (anonymous):

Please help! Solve: -2=log5x

OpenStudy (anonymous):

0.002

OpenStudy (anonymous):

Options are: x=25 x=-10 x=1/25 x=1/10

OpenStudy (anonymous):

this is log base 10?

OpenStudy (anonymous):

\[-2=\log_{5}x\]

OpenStudy (anonymous):

ok so x=5^-2 = 1/25